OpenStudy (anonymous):

Look at the formula for finding distance. d = rt Which of the following is the formula to find time (t)?

OpenStudy (anonymous):

To find t, make t the subject of the equation. Example; y = xyz Question: find z? y / xy = z

OpenStudy (anonymous):

A. t = d r B. t = d - rC. t = r d D. t = dr

OpenStudy (anonymous):

those are my options